Please solve the problem below. I would really like to see work shown so I can understand the concepts and the things I am doing incorrectly. |(1 point) A mass of 4 kg stretches a spring 40 cm. The mass is acted on by an external force of F(t) = 97 cos(0.5t) N and moFind the general solution to your differential equation from the previous part. Use C1 and C2 to denote arbitrary constants.

|(1 point) A mass of 4 kg stretches a spring 40 cm. The mass is acted on by an external force of F(t) = 97 cos(0.5t) N and moves in a medium that imparts a viscous force of 4 N when the speed of the mass is 8 cm/s. If the mass is set in motion from its equilibrium position with an initial velocity of 6 cm/s, formulate the initial value problem describing the motion of the mass. Note that the acceleration due to gravity, g, is g = 980 cm/s2
Find the general solution to your differential equation from the previous part. Use C1 and C2 to denote arbitrary constants. Use t for independent variable to represent the time elapsed in seconds. eA-(1/16) C1"sin(sqrt(6271/4)/8 t) e^-(1/16)"C2"cos(sqrt(6271/4)/8 "t) Solve the differential equation with initial conditions describing the motion/the displacement y(t) of the mass from its equilibrium position. y(t) = cm
Answer #1

Sol:- (d) Considethe Given IvP 91 Cos o.st) 4y +0 5yt 98y Wrth y co)= amd The chovacterChic fs eaton Amto m +98 =6 -o-5t C+(o-5t)) 91-O.5D Cat COS 9Co-5 (a4) tCo.5) (9S8th Cos C0-st) afain Rephu by - Co. J Cos Co. st) 48.51 2. 48.5D Cos Co.5t) + 0(e) Initial blition OSe Lohb 9409. 15 9409 าาาาา 9409- 125 and co) 6 9409 1409 t6 (9409-I5) 4.949 4.949 (4-941t)1T11 Sin 6.